# Evolution of reconnection flux during eruption of magnetic flux ropes

### v2

@@ 5 Summary and Discussion
The work was supported by Indo-US Science and Technology Forum (IUSSTF/JC-113/2019). SSM acknowledge NOVA HPC and VINO server of IIA used to perform the numerical simulation and analysis. P.C. acknowledges the DiRAC Data Intensive service (DIaL2) at the University of Leicester (project id: dp261), managed by the University of Leicester Research Computing Service on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The DiRAC service at Leicester was funded by BEIS, UKRI and STFC capital funding and STFC operations grants. DiRAC is part of the UKRI Digital Research Infrastructure. P.C. also acknowledges the allocation of computing resources at the PDC Center for High Performance Computing at KTH in Stockholm, funded by the National Academic Infrastructure for Supercomputing in Sweden. RS acknowledges support from the project EFESIS (Exploring the Formation, Evolution and Space-weather Impact of Sheath-regions), under the Academy of Finland Grant 350015. Additionally, authors thank the NASA SDO team for providing valuable HMI and AIA data. SDO is a mission under NASA’s Living with a Star program. Open access is funded by Helsinki University Library.

### v3

@@ Acknowledgment
The work was supported by Indo-US Science and Technology Forum (IUSSTF/JC-113/2019). SSM acknowledge NOVA HPC and VINO server of IIA used to perform the numerical simulation and analysis. P.C. acknowledges the DiRAC Data Intensive service (DIaL2) at the University of Leicester (project id: dp261), managed by the University of Leicester Research Computing Service and CSD3 service at the University of Cambridge on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The DiRAC service was funded by BEIS, UKRI and STFC capital funding and STFC operations grants. DiRAC is part of the UKRI Digital Research Infrastructure. P.C. also acknowledges the allocation of computing resources at the PDC Center for High Performance Computing at KTH in Stockholm, funded by the National Academic Infrastructure for Supercomputing in Sweden. RS acknowledges support from the Research Council of Finland Grant 350015. Additionally, authors thank the NASA SDO team for providing valuable HMI and AIA data. SDO is a mission under NASA’s Living with a Star program. Open access is funded by Helsinki University Library.